Determine the Q-point components of this circuit. Use APPROXIMATE analysis. Given: A voltage-divider biasing circuit with Vcc 14V, Initial Beta 80, Final Beta Determine VB. VE. lE. VCE VB at Beta of 80- VB at Beta of 120 Use 2 decimal places. 120 V, VE at Beta of 80- V, lE at Beta of 80- mA, VCE at Beta of 80- V, VE at Beta of 120 V, lE at Beta of 120 mA, VCE at Beta of 120-
Answer #1

B 14 kn FE Initial下二80 14 Bg 5 28-l 殀. Re +2.88-o. 3-11-11 +Cl+80) 1-4 2. (8 124.51TE二1.38mA 8 - O-01H 1, 38b-リ 80 = 1.93V2.88 V 120 (0-0)2) 8 Te = t-4514-C.uu(M) r (.unt.4)) Ce . B 12o